### v2

# Orchid: Flexible and Data-Dependent Convolution for Sequence Modeling

Abstract
In the rapidly evolving field of deep learning, the demand for models that are both expressive and computationally efficient has never been more critical. This paper introduces Orchid, a novel architecture designed to address the quadratic complexity of traditional attention mechanisms without compromising the ability to capture long-range dependencies and in-context learning. At the core of this architecture lies a new data-dependent global convolution layer, which contextually adapts its kernel conditioned on input sequence using a dedicated conditioning neural network. We design two simple conditioning networks that maintain shift equivariance in our data-dependent convolution operation. The dynamic nature of the proposed convolution kernel grants Orchid high expressivity while maintaining quasilinear scalability for long sequences. We evaluate the proposed model across multiple domains, including language modeling and image classification, to highlight its performance and generality. Our experiments demonstrate that this architecture not only outperforms traditional attention-based architectures such as BERT and Vision Transformers with smaller model sizes, but also extends the feasible sequence length beyond the limitations of the dense attention layers. This achievement represents a significant step towards more efficient and scalable deep learning models for sequence modeling.
1 Introduction
In modern deep neural networks, attention mechanisms have emerged as a gold standard, pivotal in domains such as natural language processing, image, and audio processing, and even complex fields like biology  (vaswani2018attention; dosovitskiy2020image; dwivedi2020generalization). However, despite their strong sequence analysis capabilities, these sequence modeling mechanisms suffer from their high computational complexity, which scales quadratically with sequence length, hindering their application to long-context tasks. This complexity has driven a shift towards innovative solutions to overcome this computational barrier, enabling analysis of long sequences in areas like genomics, DNA sequencing, and the creation of long musical compositions.
In the past years, researchers have explored various strategies to tackle the computational bottleneck of traditional dense attention layers (tay2022efficient). One key strategy involves sparsifying the dense attention matrix. Instead of calculating the entire matrix, qiu2019blockwise; parmar2018image focus on specific local blocks of the receptive fields of sequences by chunking them into fixed-size blocks. Moreover, Sparse Transformer (child2019generating), Longformer (beltagy2020longformer) and BigBird (zaheer2020big) use strided attention patterns combined with local sliding windows to reduce computation. In contrast to using pre-determined patterns, other techniques include learning to cluster/sort tokens based on a similarity function, thereby enhancing the global view of the sequence, as seen in Reformer (kitaevreformer), Routing Transformer (roy2020efficient) Sparse Sinkhorn attention (tay2020sparse). Another approach involves low-rank approximations of the self-attention matrix, leveraging the insight that these matrices often exhibit low-rank properties, as demonstrated by Linformer (wang2020linformer) which projects keys and values matrices to lower-dimensional representation matrices. Another paradigm to reduce quadratic computation cost, is to replace the dot-product similarity between keys and query matrices of attention mechanism with a kernel function and avoid explicitly computing the attention matrix (katharopoulos2020transformers). Notable examples in this family include Performers (choromanski2020rethinking), Random Feature Attention (peng2021random) that are based on random feature approximation of the kernel function. Additionally, some models leverage a combinations of such techniques to design an efficient transformer (zhu2021long; zhang2021poolingformer). However, while these methods significantly reduce computational overhead, they may sacrifice expressiveness and performance, often requiring hybrid approaches that combine them with dense attention layers (mehta2022long; fu2023hungry). On the other hand, recent works have aimed at sparsifying dense linear layers, used for feature mixing in Transformer blocks, to tackle another major source of high computation and memory demand in large models (dao2022monarch; chen2021pixelated; chen2021scatterbrain).
Finding sub-quadratic and hardware-efficient mixing operators that are also expressive remains a significant challenge. Recent studies have explored attention-free solutions, particularly using state space models (SSMs) (gu2021efficiently; mehta2022long; wang2022pretraining; fu2023hungry; orvieto2023resurrecting; gu2023mamba; de2024griffin), and long convolutions (romero2021ckconv; li2022makes; poli2023hyena). A state space model characterizes a dynamical system’s behavior in terms of its internal state using a state equation, describing the dynamics of the system using first-order differential equations over the states, and an observation equation, relating state variables to observed outputs. [^1] Notably, state space models, in general, include the recurrent layers such as RNN and LSTM (hochreiter1997LSTM) as special cases. A key insight is that, most of these SSM models can be formulated as a long convolution model between the input and output sequences (gu2021combining), allowing parallel and efficient training. However, recent work by poli2023hyena demonstrated that directly parameterizing the filter impulse response of the long-convolution leads to an even more expressive sequence mixing layer.
This paper proposes a novel data-dependent convolution mechanism to tackle the inherent quadratic complexity of traditional attention mechanisms, while maintaining the model’s ability to capture long-range dependencies and in-context learning. The data-dependent convolution layer contextually adapts its kernel based on input data using a dedicated conditioning neural network. We design two simple yet effective conditioning networks that maintain shift equivariance in the adaptive convolution operation. By combining these adaptive mechanisms with gating operations, our proposed model—named Orchid—achieves high expressivity while offering quasilinear scalability (with a complexity of $\mathcal{O}(L\log L)$) for long sequences. Evaluation across various domains, including language modeling and image classification, presented in section LABEL:sec:experiments and Appendix, demonstrates the Orchid architecture’s performance and generality, outperforming attention-based architectures, like BERT and Vision Transformers, with smaller model sizes. Moreover, its allows for handling very large sequence lengths that are beyond the limitations of the dense attention layers. This achievement lays the foundation for further advancements in more efficient and scalable sequence modeling architectures.
2 Background
Figure 2.1: Orchid block. $\odot$ and * denote element-wise multiplication and the convolution operator, respectively. The convolutions are implemented in the frequency domain using FFT. On the right two different conditioning networks, introduced in equations (2) and (3) as shift-invariant convolution kernels, are depicted. In this model, the convolution is performed efficiently in the spectral domain, so the kernel in the frequency domain, $h^{\mathcal{F}}=h_{0}^{\mathcal{F}}+h^{\mathcal{F}}_{\theta}({\bm{x}})$, is computed.
Self-Attention Mechanism:
Given a length-$L$ sequence of embeddings (tokens) ${\bm{x}}=(x_{1},x_{2},\ldots,x_{L})$, the self-attention layer generates a new sequence by computing a weighted sum of these embeddings. To achieve this, it linearly project ${\bm{x}}$ into three components: queries (${\bm{Q}}$), keys (${\bm{K}}$), and values (${\bm{V}}$), as: ${\bm{Q}}={\bm{x}}{\bm{W}}^{Q},\quad{\bm{K}}={\bm{x}}{\bm{W}}^{K},\quad{\bm{V}}={\bm{x}}{\bm{W}}^{V}.$ Each individual attention mechanism within a multi-head self-attention layer operates as a dense linear transformation, expressed as:
${\bm{y}}=\texttt{SA}({\bm{Q}},{\bm{K}},{\bm{V}})=\texttt{SoftMax}\left(\frac{{\bm{Q}}{\bm{K}}^{T}}{\sqrt{d_{k}}}\right){\bm{V}}={\bm{A}}(x){\bm{V}},$
where the matrix ${\bm{A}}(x)$ contains the normalized attention scores between each pair of tokens. This description of the attention layer highlights its notable benefits, including its capability to capture long-range dependencies using a sublinear parameter count. The attention mechanism enables direct computation of interactions between any two positions in the input sequence, regardless of their distance, without a corresponding rise in parameter counts. Additionally, the attention layer implements a data-dependent dense linear filter, which effectively filter the input based on on weights conditioned by a mapping of the data. This property makes it expressive and flexible enough to encode a large family of linear functions. However, these advantages come at the expense of quadratic computational complexity and memory costs.
This motivates us to develop an efficient and scalable data-dependent convolution mechanism, featuring an adaptive kernel that adjusts based on the input data. The kernel size of this convolution layer is as long as the input sequence length, enabling the capture of long-range dependencies across the input sequence while maintaining high scalability.
Linear Convolution:
Discrete-time linear convolution is a fundamental operation in digital signal processing that calculates the output as the weighted sum of the finite-length input ${\bm{x}}$ with shifted versions of the convolution kernel, ${\bm{h}}$, also known as the impulse response of a linear time-invariant (LTI) system.222 While generally ${\bm{x}}$ represents a sequence of D-dimensional embeddings with length $L$, all the sequence mixing operators, including data-dependent convolutions, Fourier transforms and depthwise 1D convolutions (Conv1d) are performed along the sequence dimension. Therefore, for clarity and without loss of generality, we assume $D=1$ and ${\bm{x}}~{}\in~{}\mathbb{R}^{L}$. For a full list of notation definition, please refer to Appendix LABEL:apdx:not_def. Formally, it can be written as
${\bm{y}}[t]=({\bm{h}}*{\bm{x}})[t]\triangleq\sum_{\ell=0}^{L-1}h[t-\ell]x[\ell].$
In this definition, the output is a linear filter of the zero-padded input and convolution kernel. However, other padding schemes leads to different forms of convolution. A well-known form is circular convolution, defined as
${\bm{y}}[t~{}{\sf mod}~{}L]=({\bm{h}}\circledast{\bm{x}})[t]\triangleq\sum_{\ell=0}^{L-1}h[t-\ell~{}{\sf mod}~{}L]x[\ell],$
which is equivalent to the linear convolution of two sequences if one is cyclically padded at its edges.
Global Convolution and Fast Convolution Algorithm:
Standard convolution layers, explicitly parameterized with a short kernel, struggle to capture long-range dependencies in sequential data. Extending the kernel to match the input length enables modeling such dependencies but leads to linear growth in parameter counts and quadratic computational complexity. To mitigate the parameter growth challenge, the kernel of global (a.k.a. long) convolution can be implicitly parameterized using a multilayer perceptron (MLP), a technique that has been shown to maintain sub-linear parameter scaling (karami2019invertible; romero2021ckconv; li2022makes; poli2023hyena). Furthermore, a key advantage of convolution operators is that, leveraging the convolution theorem, convolution operators can be efficiently computed in the frequency domain using Fast Fourier Transform (FFT) algorithms, thereby reducing the computational complexity to $\mathcal{O}(L\log L)$ (cooley1965FFT). Formally, the linear convolution can be expressed in the frequency domain as $\hat{{\bm{y}}}=\mathcal{F}^{-1}(\mathcal{F}(\hat{{\bm{h}}})\odot\mathcal{F}(\hat{{\bm{x}}}))={\bm{T}}^{-1}({\bm{h}}^{\mathcal{F}}\odot{\bm{T}}\hat{{\bm{x}}})$, where ${\bm{T}}$ is the DFT matrix, $\mathcal{F}$ denotes the discrete Fourier transformation, and $\hat{{\bm{x}}}$ denotes the zero-padded signal, defined as $\hat{{\bm{x}}}\triangleq\texttt{pad}_{2L}({\bm{x}})=[\mathbf{0}_{L};~{}{\bm{x}}]$. Additionally, the circular convolution can be simply computed as: ${\bm{y}}={\bm{h}}\circledast{\bm{x}}=\mathcal{F}^{-1}(\mathcal{F}({{\bm{h}}})\odot\mathcal{F}({{\bm{x}}}))$.
3 Orchid Operator
This section introduces the Data-Dependent Convolution Filter, a novel operator aimed at increasing the expressiveness of long convolution operations. This operator serves as the foundational building block for the Orchid layer, which we will explore later in the section.
3.1 Data-Dependent Convolution Filter
We hypothesize that making the convolutional kernel data-dependent allows the filter to adapt to the specific characteristics of its input, potentially capturing more complex patterns within the sequence. Formally, this input-dependent filter is defined as:
${\bm{y}}=h_{\theta}({\bm{x}})*{\bm{x}}=\text{NN}_{\theta}({\bm{x}})*{\bm{x}}$ (1)
The key innovation is to replace the static convolutional kernel with a conditionally generated one controlled by the input data. This is achieved through a conditioning network, denoted as $h_{\theta}({\bm{x}})=\text{NN}_{\theta}({\bm{x}})$, a neural network parameterized by $\theta$. The conditioning network outputs a vector matching the input sequence in length. This allows each input token to ’attend’ to the entire sequence with personalized, adaptive weights derived from its specific representation. Convolving the surrounding context using this data-dependent weighting scheme can potentially offer more effective sequence mixing compared to conventional static convolutions.
3.2 Preserving Shift-Equivariance in Data-Dependent Convolution
A fundamental property of discrete convolution is shift equivariance, meaning that shifting the input by a certain amount leads to a corresponding shift in the output (disregarding boundary effects). This is formally expressed for circular convolution as: $\texttt{shift}_{m}({\bm{y}})={\bm{h}}\circledast\texttt{shift}_{m}({\bm{x}})$ (bronstein2021geometric), where this property holds exactly regardless of boundary conditions. The shift operation is defined as $\texttt{shift}_{m}({\bm{x}})[t]\triangleq{\bm{x}}[t+m]$.
This property is particularly important because it ensures the operator’s response is robust to shift of features within the input, thereby enhancing the model’s generalization capabilities. This inductive bias is at the core of the widespread success of convolution operations (thomas1802tensor). Therefore, it is desirable to design conditioning network in the data-dependent convolution (1) to preserve shift equivariance property. To maintain this property for data-dependent convolution operations, it is sufficient to design filter kernel to be shift-invariant, i.e. $h(\texttt{shift}_{m}({\bm{x}}))=h({\bm{x}})$ (refer to Appendix LABEL:apdx:proofs for the proof). In the following, we present two conditioning network designs satisfying shift-invariance.
I) Phase Suppression for Shift Invariance:
A circular shift of a sequence ${\bm{u}}$ results in a linear phase shift of its frequency components: $\mathcal{F}\big{(}\texttt{shift}_{m}({\bm{u}})\big{)}[\omega]={\bm{u}}^{\mathcal{F}}[\omega]\cdot e^{-\frac{i2\pi}{L}\omega m}$ (oppenheim1999discrete). Given a shift-equivariant function $g(x)$ (such as a depthwise Conv1d()) (satisfying: $g(\texttt{shift}_{m}({\bm{x}}))=\texttt{shift}_{m}(g({\bm{x}}))$). Its frequency components after a spatial shift of its input maintain this phase shift:
$\mathcal{F}\big{(}g(\texttt{shift}_{m}({\bm{x}})\big{)}[\omega]=\mathcal{F}\left(g({\bm{x}})\right)[\omega]\cdot e^{-\frac{i2\pi}{L}\omega m}.$
By taking the magnitude (absolute value or squared) of these complex-valued frequency components, we effectively eliminate the phase shift, therefore, defining $h^{\mathcal{F}}({\bm{x}})=\big{|}\mathcal{F}\big{(}g({\bm{x}})\big{)}\big{|}$ satisfies shift-invariance property: $h^{\mathcal{F}}(\texttt{shift}_{m}({\bm{x}}))=h^{\mathcal{F}}({\bm{x}})$.
[…]
$h^{\mathcal{F}}_{\theta}({\bm{x}})=\texttt{Conv1d}\big{(}\big{|}\mathcal{F}\left(\texttt{Conv1d}({\bm{x}})\right)\big{|}\big{)}$ (2)
This architecture choice aims to minimize the number of parameters and computational overhead introduced by the conditioning network within the overall model.
II) Leveraging Cross-Correlation for Shift-Invariance:
An alternative approach to achieving shift-invariance involves computing the cross-correlation between two mapping of the input sequence. Let $k({\bm{x}})$ and $q({\bm{x}})$ be two shift-equivariant functions, satisfying: $k(\texttt{shift}_{m}({\bm{x}}))=\texttt{shift}_{m}(k({\bm{x}}))$ and $q(\texttt{shift}_{m}({\bm{x}}))=\texttt{shift}_{m}(q({\bm{x}}))$. We define $h({\bm{x}})$ as the cross-correlation of $k({\bm{x}})$ and $q({\bm{x}})$, given by:
$h({\bm{x}})[t]=(k({\bm{x}})\star q({\bm{x}}))[t]\triangleq\sum_{\ell=0}^{L-1}k({\bm{x}})[\ell]\cdot q({\bm{x}})[t+\ell~{}\text{ mod }L].$
This operation essentially slides $q({\bm{x}})$ over $k({\bm{x}})$ and measures their similarity at different offsets. Remarkably, the resulting cross-correlation function, $h({\bm{x}})$, is also shift invariant:
$h(\texttt{shift}_{m}({\bm{x}}))$ $=k(\texttt{shift}_{m}({\bm{x}}))\star q(\texttt{shift}_{m}({\bm{x}}))$
$=\texttt{shift}_{m}(k({\bm{x}}))\star\texttt{shift}_{m}(q({\bm{x}}))$
$=k({\bm{x}})\star q({\bm{x}})=h({\bm{x}})$
Furthermore, the convolution theorem enables efficient computation of the cross-correlation in the frequency domain: $h^{\mathcal{F}}({\bm{x}})=\mathcal{F}\left(k({\bm{x}})\star q({\bm{x}})\right)=k^{\mathcal{F}^{*}}({\bm{x}})\odot q^{\mathcal{F}}({\bm{x}})$ where $k^{\mathcal{F}^{*}}$ denotes the complex conjugate of $k^{\mathcal{F}}$ and $\odot$ represents element-wise multiplication.
Remark 3.1.
By setting $k({\bm{x}})=q({\bm{x}})=g({\bm{x}})$, we obtain $h^{\mathcal{F}}({\bm{x}})=|g^{\mathcal{F}}({{\bm{x}}})|^{2}$, This indicates that the cross-correlation approach generalizes the magnitude-based approach, demonstrating its versatility.
Similar to the previous approach, we employ separate 1D depth-wise short convolutions for both $k({\bm{x}})$ and $q({\bm{x}})$, followed by another convolution post cross-correlation in the frequency domain. As a result, the conditioning neural network is defined as
$h^{\mathcal{F}}_{\theta}({\bm{x}})=\texttt{Conv1d}\Bigl{(}\mathcal{F}^{*}\left(\texttt{Conv1d}({\bm{x}})\right)\odot\sigma\bigl{(}\mathcal{F}\left(\texttt{Conv1d}({\bm{x}})\right)\bigr{)}\Bigr{)}.$ (3)
Inspired by gating mechanism, we applied Sigmoid to one of this pair in the frequency domain.
Both conditioning functions, as defined in (2) and (3), are illustrated schematically in Figure 2.1.